OpenStudy (anonymous):

Find the sum of the first 26 terms of the following sequence: -12, -13, -14 . . .

OpenStudy (anonymous):

here a= -12 d=-13-(-12) =-1 n=26 so use S=(n/2) *( 2*a +(n-1)*d)